A tachometer generator used to measure speed of rotation of IC engine, has an ideal rating of 6 V per 1000 rpm, a range of 0-4000 rpm and accuracy of ± 0.4 % of full scale. If the output of tachometer generator id 18 V, what is the ideal value of the speed? What are the minimum & maximum values of speed.
1
Expert's answer
2020-09-24T10:24:34-0400

Range of tachometer generator : 0:4000 rpm

accuracy:  0.4% of full scale

output: 18 V

ideal rating: "\frac{6V}{1000}" rpm

Since 6V corresponds to a speed of 1000 rpm, 18 V of output would correspond to mean speed of 3000 rpm (unitary method).

Therefore, the ideal value of speed is 3000 rpm.

To calculate the minimum and maximum values of speed, we have to deduct and add the deviation to the ideal value of speed.

Deviation = 0.4 % of full scale reading, ie., 0.4 % of 4000.

Deviation = 16 rpm

Therefore, maximum speed = 3000 + 16 = 3016 rpm.

Minimum speed = 3000-16 = 2984 rpm.